OpenStudy (anonymous):

Multiply. Express the product in simplest form. 3 over 9 times 18 equals b A. 1 over 54 B. 2 C. 5 D. 6

OpenStudy (anonymous):

So \[\frac{ 3 }{ 9 }*18\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 3 }{ 9 }*\frac{ 18 }{ 1 } \to \frac{ 3*18 }{ 9*1 } \to \frac{ 54 }{ 9 }=?\] Those are the steps, now just simplify.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem :)